In MKS (SI units), mass is expressed in kilogram (Kg), length in meter (m) and time in second (s). What would be the units for the following?

a. Speed = distance/time b. Momentum = mass * speed

c. Kinetic energy = (1/2)*mass*speed2 d. Force = mass*acceleration

e. Work = Force * distance
Answer #1

a. SPEED = dist/time = meter/sec. =m/s

b. Momentum = mass* speed = mass*dist/time = kg.m/sec.

c. Kinetic energy = (1/2)*mass*speed2 = kg*m2/s2 = Joule

d. Force = mass*acceleration = mass* velocity/time = kg*m/s.s = kg*m/s2 = Newton (N)

e. Work = Force. distance = kg*m*m/sec2 = kg.m2/s2 = Joule